OpenStudy (anonymous):

ANSWERS GET MEDALS!! The force needed to keep a car from skidding on a curve varies inversely with the radius of the curve and jointly with the weight of the car and the square of the speed. If 3900 pounds of force keeps a 3510 pound car from skidding as it travels at 25 miles per hour on a curve of radius 500 feet, what force would keep the car from skidding on a curve of radius 900 feet at 45 miles per hour?

OpenStudy (anonymous):

F=(k/r)*W*S = k W S/r 3900 =k*3510*25/500 500*3900/(3510*25)=k Then substitute k in: New Force = k * 3510 *45 / 900

OpenStudy (anonymous):

so what would the final answer be..? what i got seems unreasonable

OpenStudy (anonymous):

I end up with 3900 lbs force, which seems reasonable in this case, since the speed increased, serving to increase force, and radius increased, serving to decrease it. It is not unlikely that the problem's numbers were designed specifically to yield this effect.

OpenStudy (anonymous):

yea thats not what i got sorry that seems much better